OpenStudy (anonymous):

What is the length of an arc of 72 degrees on a circle whose radius is 10????

OpenStudy (jdoe0001):

\(\large {\textit{arc's length}=s=\cfrac{r\theta\pi }{180}\quad \begin{cases} r\to radius\\ \theta\to \textit{angle in degrees} \end{cases}}\)

OpenStudy (anonymous):

so the answer will be 720*pi/180??